OpenStudy (anonymous):

There are two numbers. The second number is 1 larger than twice the first number. The sum of the reciprocals of the twon numbers is 7/10. What are the two numbers.

OpenStudy (phi):

There are two numbers. give a name to one of the numbers. example: call the first number "x" The second number is 1 larger than twice the first number. 1 larger than twice x 1 plus 2 times x 1+2x so the two numbers are: x and (1+2x) The sum of the reciprocals reciprocal of x means 1/x ("flip" it) the reciprocal of (1+2x) is 1/(1+2x) now add these reciprocals (they add up to 7/10) \[ \frac{1}{x}+\frac{1}{(2x+1)} = \frac{7}{10} \]